OpenStudy (anonymous):

what is the equation for a line tangent to x2+y2=25 at point (3,4)? and how do you get that answer?

OpenStudy (anonymous):

Take a derivative. Plug in (3,4) that gives you the slope. Then use point slope form: \[y-y_0=m(x-x_0)\] To get the equation of the line.

OpenStudy (amistre64):

y = f'(3) -3f'(3) +4

OpenStudy (amistre64):

the derivative is determined by the same rules that apply to any derivative; x^2 goes to 2x.x' ; y^2 goes to 2y.y' and 25 goes to 0 2x.x' + 2y.y' = 0 ; solve for y'

OpenStudy (anonymous):

I'm probably the worst at math. what's a derivative?

OpenStudy (amistre64):

a derivative is: \[\lim_{h->0}\frac{f(x+h)-f(x)}{h}\] but that is just how we learn to get the derivative; after that we learn that there are simple rules that derivatves follow

OpenStudy (amistre64):

the derivative of a function allows us to see the slope of the tangent line at any given point of that function

OpenStudy (amistre64):

in any event, I did the hard stuff for you; now all you have to do is algebra this equation to solve for y': 2x.x' + 2y.y' = 0 ; solve for y'

OpenStudy (anonymous):

okay thank you so much! this will really help with my placement test coming up! i hate math!